# PSZ2 G181.06+48.47 I: X-ray exploration of a low-mass cluster with exceptionally-distant radio relics

### v1

@@ 4.2 Global properties and cluster mass estimation
The global temperature obtained by jointly fitting XMM-Newton MOS and PN data is $kT_{500}=3.86^{+0.27}_{-0.29}$ keV (see Table 2). The global luminosities within $r_{\rm 500,SZ}$ calculated within a range of energy bands are also summarized in Table 2. PSZ2 G181.06+48.47 has comparatively a lower temperature and luminosity than samples of clusters studied with XMM-Newton observations (e.g. Pratt et al., 2009; Lovisari et al., 2020).
@@
Multiple $M_{500}-kT_{\rm X}$ scaling relations have been presented in the literature, derived from different cluster samples. For example, Lovisari et al. (2020) considered the effect non-cool core clusters have on scaling relations and derived scaling relations using a representative sample of 120 Planck clusters, which included both relaxed and disturbed clusters. The Planck selection and the XMM-Newton-derived X-ray properties should provide a good comparison to our cluster. The presence of two subclusters introduces uncertainty in excising any potential cool core, so we use the relation derived using global, core-included temperatures. Adopting the relation for merging clusters, as appropriate for the highly disturbed PSZ2 G181.06+48.47 cluster, we obtain a mass of $M_{500,X}=2.57^{+0.37}_{-0.38}\times 10^{14}\;\mathrm{M_{\odot}}$. Using the Lovisari et al. (2020) relation for the entire cluster sample results in a slightly higher, but consistent value ($2.71^{+0.32}_{-0.34}\times 10^{14}\;\mathrm{M_{\odot}}$). The resulting $r_{\rm 500,X}$ estimated from the disturbed cluster scaling relation is $0.90\pm 0.05\,\mathrm{Mpc}$.
@@
The X-ray derived mass is lower than the Planck mass ($M_{\rm 500,SZ}=(4.2\pm 0.5)\times 10^{14}\;{\rm M_{\odot}}$), a difference which is significant at the $2.6\sigma$ level. The discrepancy is likely to be even higher, since shock heating can boost the temperature and bias-high temperature-derived masses (as found in A 2108 by Schellenberger et al., 2022). Planck Collaboration et al. (2016) caution that the uncertainties provided represent a lower limit on the real uncertainty, as they do not include all known sources of errors. Another contribution to the discrepancy could be the different centers used in the two measurements (e.g. the midpoint between the subclusters was used for the X-ray measurement, while the Planck measurement is centered closer to the main peak in the south) or the inclusion in the mass estimation of the SZ signal from any of the background clusters identified in the field. As expected, $r_{\rm 500,X}$ is also significantly lower than the Planck value of $r_{\rm 500,SZ}=1.06\pm 0.05\,\mathrm{Mpc}$ at the $2.2\sigma$ level.
@@
Ahn et al. (in prep) derive a WL mass for the system of $M_{\rm 500,WL}=2.90^{+0.75}_{-0.69}\times 10^{14}\;\mathrm{M_{\odot}}$ by fitting two superpositioned Navarro-Frenk-White halos (Navarro et al., 1997). The WL mass is in excellent agreement with the X-ray derived mass. The X-ray and WL-derived masses place the PSZ2 G181.06+48.47 among the lowest mass clusters to host double radio relics. Out of the 30 known double radio relic systems (see Section 6 for details), only seven have masses $\lesssim 3\times 10^{14}\,{\rm M_{\odot}}$.
@@
SDSS DR18 spectroscopy unveils an overdensity of galaxies at $z\sim 0.24$ in the vicinity of the cluster (see Figure 6). By modeling the limited spectroscopy (17 galaxies) within $r_{\rm 500,SZ}$ of the cluster center with a Gaussian distribution, we estimate the cluster redshift as $z=0.2335\pm 0.0035$, with a velocity dispersion of $\sim 1290$ km s-1 and an equivalent dynamical mass of $2.5\times 10^{15}$ M (following scaling relations from Saro et al., 2013). Li et al. (2023) have shown that dynamical masses for unrelaxed clusters can be approximately 50% higher than those derived from X-ray data. Given the disturbed nature of the cluster, this mass estimation is unsurprisingly higher than the cluster mass estimated from the gas properties (X-ray temperature and SZ signal). The spectroscopic velocities of the two BCGs ($z_{\rm N}=0.234759\pm 0.000046$ and $z_{\rm S}=0.231484\pm 0.000036$) are consistent within the velocity distribution calculated for the cluster. However, the line-of-sight (LOS) velocity difference might indicate that the subclusters are moving towards or away from one another at 982 km s-1 along the LOS, consistent with an ongoing merger with a LOS component. The high LOS difference between the two BCGs in PSZ2 G181.06+48.47 could further inflate the dynamical mass estimate, as was found in other clusters such as Abell 115 (Kim et al., 2019).
@@ 4.3 Properties of the subclusters
Note that while the total mass measured by the WL method (Ahn et al., in prep) is in excellent agreement with the X-ray-derived mass, the mass ratios between the subclusters slightly differ. Both the temperature and luminosity ratios of the two subclusters suggest that the southern subcluster is about 1.2–1.4 times more massive than the northern subcluster, while the WL finds a mass ratio that is closer to 1:3 ($3.0\pm 1.3$). On the one side, the proximity between the two subclumps might result in some projection effects that slightly affect the temperature estimate. On the other side, the WL mass map reveals extensions towards the southeast and northwest of the southern subcluster core, but no corresponding feature is found in the X-ray map, possibly indicating subgroups that were stripped of gas during their infall into the PSZ2 G181.06+48.47 system (Eckert et al., 2014; Ebeling et al., 2017).
Figure 7: Temperature, density, pressure, and entropy maps focusing on the central $7\arcmin\times 7\arcmin$ are of PSZ2 G181.06+48.47. X-ray surface brightness contours are drawn in gray at $[4,8,12,16]\times 10^{-8}\,{\rm cts/s}$ and the edges are highlighted with black dashed curves. The temperature is shown in a linear scale, while the other maps are shown with a square-root stretch.
@@ 4.4 Thermodynamic properties
The highest densities coincide with the location of the subclusters, as implied by the X-ray map and the positions of the BCGs. The density map (upper right panel in Figure 7) reveals a high density region in the cores of the two subclusters and the connecting bridge, flanked by areas of lower density. Outside of the cores, the density drops more sharply along the merger axis (NE of the NE core and SW of the SW core), than in the perpendicular direction. The density might drop by about $30-40$% outside of the central region. Pressure peaks correspond to both subcluster cores, with the southern core possibly showing peak pressures $\sim 60-70$% above the northern core (see the lower left panel in Figure 7).
The low-entropy southern core is embedded in a comparatively higher entropy region than the northern core (lower right panel in Figure 7). In addition, a bridge of low entropy gas connects the two subclusters, with the lowest entropies found in the northern core. Note however that the relative BCG velocities suggest a LOS component to the merger, which could imply that the connecting bridge also has a LOS component. If all of that gas has the same temperature, the measured temperature would not be affected, but the normalization and inferred density will be driven up. The combined effect would artificially raise the projected pseudo-entropy. Simulations have shown that, in the absence of non-gravitational effects, the low-entropy, dense, cold gas settles to the center of relaxed galaxy clusters (e.g. Voit et al., 2002; McCarthy et al., 2008; Power et al., 2014). The low-entropy gas corresponding to each of the cores indicates that the two subclusters might have been cool-core, relaxed clusters before the merger event. While the merger likely did not fully disrupt either of the two subcluster cores, there is stronger evidence for disruption in the southern core. Towards the west, perpendicular to the merger axis, regions of possibly hot ($\sim$5.1 keV) and cold ($\sim 2.9$ keV) plasma are located next to each other. The cold area has higher densities, lower pressures, and significantly lower entropies. Such low entropy plumes can be formed most cleanly shortly ($\sim 500$ Myr) after head-on collisions between clusters of similar masses (ZuHone, 2011). The features observed here could be related to the southeast-northwest extension observed in the WL mass map (Ahn et al., in prep), possibly indicating a third-body collision experienced by the southern subcluster region, or sloshing activity potentially triggered by the core passage.
@@ 5.2 Discontinuities at the radio relic positions
As shown in Figure 9, while no clear density jumps were detected in these regions, we can use a broken-power law model to place limits on the X-ray Mach number of the shocks. The position of the shocks was fixed to the outer edge of the radio relic, where the putative location of the shock is expected to fall (see Figure 3). For the NE relic, from the fit compression ratio of $1.00\pm 0.13$, we place a $5\sigma$ upper limit on the Mach number of $\mathcal{M}_{\rm X,\rho}<1.43$, while for the SW relic the compression ratio $1.00\pm 0.17$ yields a $5\sigma$ limit of $\mathcal{M}_{\rm X,\rho}<1.57$. Note that if there are significant projection effects at play caused by the merger being offset from the place-of-the-sky, the intrinsic density-derived Mach number might be higher (e.g. Ohmura et al., 2023). We further consider the potential uncertainty in the X-ray derived Mach numbers due to the assumed ratio of specific heats $\gamma=5/3$. For fully relativistic gas, the specific heat ratio can approach $4/3$, resulting in $\mathcal{M}_{\rm X}<1.38$ and $<1.50$ for the NE and SW shocks, respectively, which suggests the shocks may be even weaker. Since the extraction regions for the relics are located partially beyond $r_{\rm 500,P}$, the low S/N regime impacts our ability to make temperature measurements. We are, therefore, unable to constrain the temperatures in the pre- and post-shock regions at the relic locations.
@@
As described in Rajpurohit et al. (submitted), a significant spectral gradient has been observed for the SW relic in the spectral index map, which rules out its classification as a double-lobed radio galaxy, and indicates a radio relic nature. The integrated spectral index implies a high Mach number shock $\mathcal{M}_{\rm R}=4.8^{+2.3}_{-0.9}$ under the assumption of diffusive shock acceleration (DSA), significantly above the upper limit for the X-ray Mach number $\mathcal{M}_{\rm X,\rho}<1.57$, further supporting a scenario in which the merger has a significant LOS component.
@@
Radio Mach numbers typically lie above the corresponding X-ray Mach numbers (e.g., van Weeren et al. 2016; Hoang et al. 2018; Stuardi et al. 2019 and van Weeren et al. 2019 for a review). From the X-ray perspective, the estimated Mach number could be lowered by a number of observational factors, for example, the mixing of plasma with different properties within the large sectors required by the low number counts at the cluster periphery and projection effects (e.g. Ohmura et al., 2023). Further, the merger axis of the cluster is likely tilted at an angle $\geq 45\degree$ with respect to the plane of the sky (e.g. velocity difference between BCGs and radio considerations from Rajpurohit et al., submitted), which could result in unaccounted-for projection effects, such as mixing of plasma along the LOS, as well as a “blurring” of the X-ray discontinuity. As further discussed in Rajpurohit et al. (submitted), the discrepancy is most likely explained by the intrinsic differences in which part of the Mach number distribution the X-ray and radio trace, with X-ray observations probing the average of the Mach number distribution, while the radio probes the high Mach number tail (e.g. Wittor et al., 2021; Dominguez-Fernandez et al., 2020).
@@ 6 Comparison with other double radio relic systems
Figure 11: Radio power at 150 MHz versus $M_{500}$ relation for double relic clusters. We scaled the 1.4 GHz radio power to 150 MHz using the spectral indices of each relic. Moreover, 14 relics without spectral information are included, but they do not significantly affect the fitting results. The orthogonal regression fit is shown as a dashed line.
de Gasperin et al. (2014) observed a positive correlation between the 1.4 GHz radio power of double relics, the cluster mass, and relic LLS. To perform a comprehensive analysis of the scaling relations, we systematically searched the literature for all double radio relics identified to date. Double relics are thought to trace mergers in the plane of the sky, thus making them less affected by projection effects and ideal for studying scaling relations. In Table LABEL:table:double_radio_relic, we compile data for 30 galaxy clusters hosting double relics, including a total of 60 relics. This sample includes 12 new systems containing double radio relics compared to Jones et al. (2023) and de Gasperin et al. (2014). The clusters span a mass range of $1.7\times 10^{14}M_{\odot}<M_{500}<14\times 10^{14}M_{\odot}$ and a redshift range of $0.870.04<z<$ 0.87. We estimated radio power as:
@@ 6.1 Double relic scaling relations
The top left panel of Figure 10 reveals a strong correlation ($p=6\times 10^{-8}$) between double relic power and cluster mass, providing compelling evidence to reject the null hypothesis. A discernible trend suggests that high-redshift clusters tend to be more massive, while relics with lower radio power are predominantly found in nearby clusters. The median redshift of our sample is 0.233, so about half the sample is located above this redshift. Since seven of the 10 clusters at $z>0.3$ were discovered by Planck and the other three by X-ray observations, the mass limit evolves with redshift, naturally resulting in a bias toward massive clusters at higher redshifts (e.g. Planck Collaboration et al., 2016). Despite the slightly different cosmology and the expanded sample which doubles the size of our sample compared to de Gasperin et al. (2014) and improves upon it with updated values for radio power, spectral index, and related properties, the slope of the radio power at 1.4 GHz vs. cluster mass relation is consistent within the errors: our value of $3.10\pm 0.59$ compared to their value of $2.83\pm 0.39$. When the 14 relics without spectral index measurements are excluded, we obtain a slope of $2.78\pm 0.58$. The radio power of relics is theoretically expected to scale with the cluster mass following the relation $P\propto M^{5/3}$ (de Gasperin et al., 2014). However, both our and the de Gasperin et al. (2014) observations indicate a different mass dependency, with $P_{1.4\,\mathrm{GHz}}\propto M^{3.10\pm 0.59}$, which is steeper than expected theoretically (i.e, $M^{5/3}$), at the $2.4\sigma$ level. de Gasperin et al. (2014) suggest that this discrepancy arises from parameters that depend on cluster mass which, in turn, drive particle acceleration efficiencies or magnetic field strengths. It is worth noting that Jones et al. (2023), using their sample and that of de Gasperin et al. (2014), reported a much steeper dependency of $P_{150\,\mathrm{MHz}}\propto M^{5.2\pm 1.2}$ indicating a significant difference between the relations at 150 MHz and 1.4 GHz. To investigate this, we also compare the power-mass scaling relations of double relics in our sample with those of Jones et al. (2023) at 150 MHz. To do this, we converted the 1.4 GHz radio power to 150 MHz using the spectral index. For the 14 relics without spectral index information, we assumed a constant value of $\alpha=-1.3$. The resulting plot is shown in Figure 11. In contrast to Jones et al. (2023), we find that the mass dependence on both 150 MHz and 1.4 GHz radio power is strikingly similar. Specifically, the measured correlation slope is $3.31\pm 0.59$. We emphasize that, when excluding relics lacking spectral index information, the measured slope remains consistent, with a value of $3.19\pm 0.61$.
@@
We find significant correlations between LLS vs. $D_{\rm RR}$ as well as $D_{\rm RC}$, see Figure 10 middle panel. The shock surfaces could thus increase in size with greater distance from the cluster center, resulting in larger relics farther from the cluster center, consistent with prior findings by Bonafede et al. (2012), de Gasperin et al. (2014) and Jones et al. (2023). Our best fit slopes are $\mathrm{LLS}\propto D_{\rm RC}^{1.16\pm 0.35}$ and $\mathrm{LLS}\propto D_{\rm RR}^{1.38\pm 0.48}$. The latter slope is more or less consistent de Gasperin et al. (2014), namely $\mathrm{LLS}\propto D_{\rm RR}^{2.13\pm 0.25}$. Conversely, the slope for LLS vs. $D_{\rm RC}$ is consistent with the results of Jones et al. (2023) at 150 MHz, supporting their conclusion that LLS does not vary significantly at lower frequencies.

### v3

@@ 4.2 Global properties and cluster mass estimation
The global temperature obtained by jointly fitting XMM-Newton MOS and PN data is $kT_{500}=3.62^{+0.15}_{-0.07}$ keV (see Table 2). The global luminosities within $r_{\rm 500,SZ}$ calculated within a range of energy bands are also summarized in Table 2. PSZ2 G181.06+48.47 has comparatively a lower temperature and luminosity than samples of clusters studied with XMM-Newton observations (e.g. Pratt et al., 2009; Lovisari et al., 2020).
@@
Multiple $M_{500}-kT_{\rm X}$ scaling relations have been presented in the literature, derived from different cluster samples. For example, Lovisari et al. (2020) considered the effect non-cool core clusters have on scaling relations and derived scaling relations using a representative sample of 120 Planck clusters, which included both relaxed and disturbed clusters. The Planck selection and the XMM-Newton-derived X-ray properties should provide a good comparison to our cluster. The presence of two subclusters introduces uncertainty in excising any potential cool core, so we use the relation derived using global, core-included temperatures. Adopting the relation for merging clusters, as appropriate for the highly disturbed PSZ2 G181.06+48.47 cluster, we obtain a mass of $M_{500,X}=2.32^{+0.29}_{-0.25}\times 10^{14}\;\mathrm{M_{\odot}}$. Using the Lovisari et al. (2020) relation for the entire cluster sample results in a slightly higher, but consistent value ($2.46^{+0.22}_{-0.16}\times 10^{14}\;\mathrm{M_{\odot}}$). The resulting $r_{\rm 500,X}$ estimated from the disturbed cluster scaling relation is $0.87\pm 0.04\,\mathrm{Mpc}$.
@@
The X-ray derived mass is lower than the Planck mass ($M_{\rm 500,SZ}=(4.2\pm 0.5)\times 10^{14}\;{\rm M_{\odot}}$), a difference which is significant at the $3.3\sigma$ level. The discrepancy is likely to be even higher, since shock heating can boost the temperature and bias-high temperature-derived masses (as found in A 2108 by Schellenberger et al., 2022). Planck Collaboration et al. (2016) cautions that the uncertainties provided represent a lower limit on the real uncertainty, as they do not include all known sources of errors. Another contribution to the discrepancy could be the different centers used in the two measurements (e.g. the midpoint between the subclusters was used for the X-ray measurement, while the Planck measurement is centered closer to the main peak in the south) or the inclusion in the mass estimation of the SZ signal from any of the background clusters identified in the field. As expected, $r_{\rm 500,X}$ is also significantly lower than the Planck value of $r_{\rm 500,SZ}=1.06\pm 0.05\,\mathrm{Mpc}$ at the $3.0\sigma$ level.
@@
Ahn et al. (2025) derive a WL mass for the system of $M_{\rm 500,WL}=2.90^{+0.75}_{-0.69}\times 10^{14}\;\mathrm{M_{\odot}}$ by fitting two superpositioned Navarro-Frenk-White halos (Navarro et al., 1997). The WL mass is in excellent agreement with the X-ray derived mass. The X-ray and WL-derived masses place the PSZ2 G181.06+48.47 among the lowest mass clusters to host double radio relics. Out of the 30 known double radio relic systems (see Section 6 for details), only seven have masses $\lesssim 3\times 10^{14}\,{\rm M_{\odot}}$.
@@
SDSS DR18 spectroscopy unveils an overdensity of galaxies at $z\sim 0.24$ in the vicinity of the cluster (see Figure 6). By modeling the limited spectroscopy (17 galaxies) within $r_{\rm 500,SZ}$ of the cluster center with a Gaussian distribution, we estimate the cluster redshift as $z=0.2335\pm 0.0035$, with a velocity dispersion of $\sim 1290$ km s-1 and an equivalent dynamical mass of $2.5\times 10^{15}$ M (following scaling relations from Saro et al., 2013). Li et al. (2023) have shown that dynamical masses for unrelaxed clusters can be approximately 50% higher than those derived from X-ray data. Given the disturbed nature of the cluster, this mass estimation is higher than the cluster mass estimated from the gas properties (X-ray temperature and SZ signal). The spectroscopic velocities of the two BCGs ($z_{\rm N}=0.234759\pm 0.000046$ and $z_{\rm S}=0.231484\pm 0.000036$) are consistent within the velocity distribution calculated for the cluster. However, the line-of-sight (LOS) velocity difference might indicate that the subclusters are moving towards or away from one another at 982 km s-1 along the LOS, consistent with an ongoing merger with a LOS component. The high LOS difference between the two BCGs in PSZ2 G181.06+48.47 could further inflate the dynamical mass estimate, as was found in other clusters such as Abell 115 (Kim et al., 2019).
@@ 4.3 Properties of the subclusters
Note that while the total mass measured by the WL method (Ahn et al., 2025) is in excellent agreement with the X-ray-derived mass, the mass ratios between the subclusters slightly differ. Both the temperature and luminosity ratios of the two subclusters suggest that the southern subcluster is about 1.2–1.4 times more massive than the northern subcluster, while the WL finds a mass ratio that is closer to 1:3 ($3.0\pm 1.3$). On the one side, the proximity between the two subclumps might result in some projection effects that slightly affect the temperature estimate. On the other side, the WL mass map reveals extensions towards the southeast and northwest of the southern subcluster core, but no corresponding feature is found in the X-ray map, possibly indicating subgroups that were stripped of gas during their infall into the PSZ2 G181.06+48.47 system (Eckert et al., 2014; Ebeling et al., 2017).
Figure 7: Temperature, density, pressure, and entropy maps focusing on the central $7\arcmin\times 7\arcmin$ area of PSZ2 G181.06+48.47. X-ray surface brightness contours are drawn in gray at $[4,8,12,16]\times 10^{-8}\,{\rm cts/s}$ and the edges are highlighted with black dashed curves. The temperature is shown in a linear scale, while the other maps are shown with a square-root stretch.
@@ 4.4 Thermodynamic properties
The highest densities coincide with the location of the subclusters, as implied by the X-ray map and the positions of the BCGs. The density map (upper right panel in Figure 7) reveals a high density region in the cores of the two subclusters and the connecting bridge, flanked by areas of lower density. Outside of the cores, the density drops more sharply along the merger axis (NE of the NE core and SW of the SW core), than in the perpendicular direction. Pressure peaks correspond to both subcluster cores, with the southern core possibly showing peak pressures $\sim 60-70$% above the northern core (see the lower left panel in Figure 7).
The low-entropy southern core is embedded in a comparatively higher entropy region than the northern core (lower right panel in Figure 7). In addition, a bridge of low entropy gas connects the two subclusters, with the lowest entropies found in the northern core. Note however that the relative BCG velocities suggest a LOS component to the merger, which could imply that the connecting bridge also has a LOS component. If all of that gas has the same temperature, the measured temperature would not be affected, but the normalization and inferred density will be driven up. The combined effect would artificially raise the projected pseudo-entropy. Simulations have shown that, in the absence of non-gravitational effects, the low-entropy, dense, cold gas settles to the center of relaxed galaxy clusters (e.g. Voit et al., 2002; McCarthy et al., 2008; Power et al., 2014). The low-entropy gas corresponding to each of the cores indicates that the two subclusters might have been cool-core, relaxed clusters before the merger event. While the merger likely did not fully disrupt either of the two subcluster cores, there is stronger evidence for disruption in the southern core. Towards the west, perpendicular to the merger axis, regions of possibly hot ($\sim 5.1$ keV) and cold ($\sim 2.9$ keV) plasma are located next to each other. The cold area has higher densities, lower pressures, and significantly lower entropies. Such low entropy plumes can be formed most cleanly shortly ($\sim 500$ Myr) after head-on collisions between clusters of similar masses (ZuHone, 2011). The features observed here could be related to the southeast-northwest extension observed in the WL mass map (Ahn et al., 2025), possibly indicating a third-body collision experienced by the southern subcluster region, or sloshing activity potentially triggered by the core passage.
@@ 5.2 Discontinuities at the radio relic positions
As shown in Figure 9, while no clear density jumps were detected in these regions, we can use a broken-power law model to place limits on the X-ray Mach number of the shocks. The position of the shocks was fixed to the outer edge of the radio relic, where the putative location of the shock is expected to fall (see Figure 4). For the NE relic, from the fit compression ratio of $1.00\pm 0.13$, we place a $5\sigma$ upper limit on the Mach number of $\mathcal{M}_{\rm X,\rho}<1.43$, while for the SW relic the compression ratio $1.00\pm 0.17$ yields a $5\sigma$ limit of $\mathcal{M}_{\rm X,\rho}<1.57$. Note that if there are significant projection effects at play caused by the merger being offset from the plane-of-the-sky, the intrinsic density-derived Mach number might be higher (e.g. Ohmura et al., 2023). We further consider the potential uncertainty in the X-ray derived Mach numbers due to the assumed ratio of specific heats $\gamma=5/3$. For fully relativistic gas, the specific heat ratio can approach $4/3$, resulting in $\mathcal{M}_{\rm X}<1.38$ and $<1.50$ for the NE and SW shocks, respectively, which suggests the shocks may be even weaker. Since the extraction regions for the relics are located partially beyond $r_{\rm 500,P}$, the low S/N regime impacts our ability to make temperature measurements. We are, therefore, unable to constrain the temperatures in the pre- and post-shock regions at the relic locations.
@@
As described in Rajpurohit et al. (2025), a significant spectral gradient has been observed for the SW relic in the spectral index map, which rules out its classification as a double-lobed radio galaxy, and indicates a radio relic nature. The integrated spectral index implies a high Mach number shock $\mathcal{M}_{\rm R}=4.8^{+2.3}_{-0.9}$ under the assumption of diffusive shock acceleration (DSA), significantly above the upper limit for the X-ray Mach number $\mathcal{M}_{\rm X,\rho}<1.57$.
@@
Radio Mach numbers typically lie above the corresponding X-ray Mach numbers (e.g., van Weeren et al. 2016; Hoang et al. 2018; Stuardi et al. 2019 and van Weeren et al. 2019 for a review). From the X-ray perspective, the estimated Mach number could be lowered by a number of observational factors, for example, the mixing of plasma with different properties within the large sectors required by the low number counts at the cluster periphery and projection effects (e.g. Ohmura et al., 2023). Further, the merger axis of the cluster is likely tilted at an angle $\geq 45\degree$ with respect to the plane-of-the-sky (e.g. velocity difference between BCGs and radio considerations from Rajpurohit et al., 2025), which could result in unaccounted-for projection effects, such as mixing of plasma along the LOS, as well as a “blurring” of the X-ray discontinuity. As further discussed in Rajpurohit et al. (2025), the discrepancy is most likely explained by the intrinsic differences in which part of the Mach number distribution the X-ray and radio trace, with X-ray observations probing the average of the Mach number distribution, while the radio probes the high Mach number tail (e.g. Wittor et al., 2021; Dominguez-Fernandez et al., 2020).
@@ 6 Comparison with other double radio relic systems
de Gasperin et al. (2014) observed a positive correlation between the 1.4 GHz radio power of double relics, the cluster mass, and relic LLS. To perform a comprehensive analysis of the scaling relations, we systematically searched the literature for all double radio relics identified to date. Double relics are thought to trace mergers in the plane-of-the-sky, thus making them less affected by projection effects and ideal for studying scaling relations (e.g. Golovich et al., 2019; van Weeren et al., 2019). In Table LABEL:table:double_radio_relic, we compile data for 30 galaxy clusters hosting double relics, including a total of 60 relics. This sample includes 12 new systems containing double radio relics compared to Jones et al. (2023) and de Gasperin et al. (2014). The clusters span a mass range of $1.7\times 10^{14}M_{\odot}<M_{500}<14\times 10^{14}M_{\odot}$ and a redshift range of $0.870.04<z<$ 0.87. We estimated radio power as:
@@ 6.1 Double relic scaling relations
Figure 10 reveals a strong correlation ($p=6\times 10^{-8}$) between double relic 1.4 GHz power and cluster mass, providing compelling evidence to reject the null hypothesis. A discernible trend suggests that high-redshift clusters tend to be more massive, while relics with lower radio power are predominantly found in nearby clusters. The median redshift of our sample is 0.233, so about half the sample is located above this redshift. Since seven of the 10 clusters at $z>0.3$ were discovered by Planck and the other three by X-ray observations, the mass limit evolves with redshift, naturally resulting in a bias toward massive clusters at higher redshifts (e.g. Planck Collaboration et al., 2016). Despite the slightly different cosmology and the expanded sample which doubles the size of our sample compared to de Gasperin et al. (2014) and improves upon it with updated values for radio power, spectral index, and related properties, the slope of the radio power at 1.4 GHz vs. cluster mass relation is consistent within the errors: our value of $3.10\pm 0.59$ compared to their value of $2.83\pm 0.39$. When the 14 relics without spectral index measurements are excluded, we obtain a slope of $2.78\pm 0.58$. The radio power of relics is theoretically expected to scale with the cluster mass following the relation $P\propto M^{5/3}$ (de Gasperin et al., 2014). However, both our and the de Gasperin et al. (2014) observations indicate a different mass dependency, with $P_{1.4\,\mathrm{GHz}}\propto M^{3.10\pm 0.59}$, which is steeper than expected theoretically (i.e, $M^{5/3}$), at the $2.4\sigma$ level. de Gasperin et al. (2014) suggest that this discrepancy arises from parameters that depend on cluster mass which, in turn, drive particle acceleration efficiencies or magnetic field strengths. It is worth noting that Jones et al. (2023), using their sample and that of de Gasperin et al. (2014), reported a much steeper dependency of $P_{150\,\mathrm{MHz}}\propto M^{5.2\pm 1.2}$ indicating a significant difference between the relations at 150 MHz and 1.4 GHz. To investigate this, we also compare the power-mass scaling relations of double relics in our sample with those of Jones et al. (2023) at 150 MHz. To do this, we converted the 1.4 GHz radio power to 150 MHz using the spectral index. For the 14 relics without spectral index information, we assumed a constant value of $\alpha=-1.3$ (top, left panel in Figure 10). In contrast to Jones et al. (2023), we find that the mass dependence on both 150 MHz and 1.4 GHz radio power is strikingly similar. Specifically, the measured correlation slope is $3.31\pm 0.59$. We emphasize that, when excluding relics lacking spectral index information, the measured slope remains consistent, with a value of $3.19\pm 0.61$.
@@
We find significant correlations between LLS vs. $D_{\rm RR}$ as well as $D_{\rm RC}$ (see Figure 10). The shock surfaces could thus increase in size with greater distance from the cluster center, resulting in larger relics farther from the cluster center, consistent with prior findings by Bonafede et al. (2012), de Gasperin et al. (2014) and Jones et al. (2023). Our best fit slopes are $\mathrm{LLS}\propto D_{\rm RC}^{1.16\pm 0.35}$ and $\mathrm{LLS}\propto D_{\rm RR}^{1.38\pm 0.48}$. The latter slope is more or less consistent with de Gasperin et al. (2014), namely $\mathrm{LLS}\propto D_{\rm RR}^{2.13\pm 0.25}$. Conversely, the slope for LLS vs. $D_{\rm RC}$ is consistent with the results of Jones et al. (2023) at 150 MHz, supporting their conclusion that LLS does not vary significantly at lower frequencies.